Find the result when 28 is decreased by 25%. Show work..
Answer by jim_thompson5910(35256)

25% of 28 = 0.25*28 = 7

So 25% of 28 = 7

So when you decrease 28 by 25%, you are subtracting 7 from 28

This means 28 is decreased by 25% = 28 - 7 = 21